OpenStudy (anonymous):

What is the slope of the line that passes through the points (2, 3) and (–2, 6)? A. 3/4 B. -3/4 C. -3/2 D. 0

OpenStudy (eric_d):

B

OpenStudy (eric_d):

6-3/-2-2=

OpenStudy (eric_d):

=B

OpenStudy (anonymous):

slope of the line= (y1-y2)/(x1-x2) here, slope of the line=(3-6)/(2+2)=-3/4

OpenStudy (eric_d):

slope is gradient

OpenStudy (eric_d):

 Always remember this formula... to fing slope/gradient,m